OpenStudy (madmerc):

What is the simplified form of ^6 sqrt x times ^6 sqrt x times ^6 sqrt x times ^6 sqrt x? x^2/3 x^3/2 4 ^6 sqrt x ^6 sqrt 4x

OpenStudy (madmerc):

\[\sqrt[6]{x} \times \sqrt[6]{x} \times \sqrt[6]{x} \times \sqrt[6]{x}\] it looks like this

OpenStudy (anonymous):

\[= \huge (\sqrt[6]{x})^4=(x)^{4/6}\]Now just reduce the fraction in the exponent X)

OpenStudy (madmerc):

so x^2/3 ?

OpenStudy (anonymous):

Yep!!

OpenStudy (madmerc):

thank you!

OpenStudy (anonymous):

You're welcome :)